# Robust and Fine-Grained Detection of AI Generated Texts

### v1

@@ 2 Related Works
A major portion of current research in detecting machine-generated content focuses on longer-form writing through binary classification. However, AI-generated misinformation is more likely to cause harm than its use in academia, making the distinction between AI and human-generated texts on social media platforms a critical challenge. Existing methods often struggle with accurately identifying AI-generated content over shorter texts. Moreover, binary classification approaches, which categorize texts as either human or AI-generated (Wang et al., 2024a), (Wang et al., 2024b), (Bhattacharjee et al., 2023), (Zellers et al., 2019b), (Macko et al., 2023), (Ghosal et al., 2023), (Dugan et al., 2024) are less practical in settings where texts could be co-authored by both humans and LLMs. In contrast, binary classification may be more effective for shorter texts commonly found on reviews and social media platforms (Macko et al., 2024a), (Ignat et al., 2024), where content typically consists of one or two sentences. Additionally, some detection works rely on detecting watermarks from AI-generated texts, (Chang et al., 2024), (Dathathri et al., 2024), (Sadasivan et al., 2024), (Zhao et al., 2023) but not all generators utilize watermarking limiting the applicability of such approaches. Few other approaches utilize statistical methods (Mitchell et al., 2023), (Kumarage et al., 2023), (Gehrmann et al., 2019b), (Hans et al., 2024), (Bao et al., 2023), but they can be prone to mis-classification against adversarial methods like rephrasing and humanizing. (Abassy et al., 2024) introduced a 4-way classification as entirely human authored, entirely llm authored, human-edited and llm-authored or llm-edited human-authored. An ideal detection system should be capable of identifying AI-generated content from any generator without depending on watermarking, especially since watermarking techniques may not be effective for shorter texts. Further an ideal detector should be robust against adversarial methods. To properly deal with co-authored text cases, a token classification approach to detect boundaries (Dugan et al., 2022), (Macko et al., 2024b) between machine authored and human authored portions might be more appropriate. Further in cases of AI usage in scenarios like academic cases, users are likely to use a proprietary LLM with better knowledge cutoffs than an open source LLM. Similarly, for AI misuse over social platforms, users are more likely to use a open-sourced model due to better flexibility and privacy. Hence, building models and benchmarks with a appropriate set of LLMs might be more applicable for practical scenarios. Many proprietary systems struggle at the task of fine-grained detection, further a large enough dataset to cover all POS-tag bi-grams of the text boundaries is required for such fine-grained detectors to work well (Kadiyala, 2024). Previous works in the similar direction include (Lee et al., 2022), (Zhang et al., 2024), (Dugan et al., 2023), (Macko et al., 2024b), (Liang et al., 2024) which utilize a dataset of limited size and limited number of generators or those less likely to be used, which might not be enough for a detector to work well on unseen domains and generators’ texts.
@@ 3 Dataset
Our dataset consists of around 2.45M samples. We used 12 different LLMs out of which 9 are popular proprietary LLMs : GPT-o1 (OpenAI, 2024), GPT-4o (etal., 2024), Gemini-1.5-Pro (DeepMind, 2024), Gemini-1.5-Flash, Claude-3.5-Sonnet (Anthropic, 2023), Claude-3.5-Haiku, Perplexity-Sonar-Large (Perplexity, 2023), Amazon-Nova-Pro (Intelligence, 2024), Amazon-Nova-Lite. We also included 3 open-source LLMs i.e Aya-23 (Aryabumi et al., 2024), Command-R-Plus (Cohere For AI, 2024), Mistral-large-2411 (Mistral AI, 2024) which produced outputs that are relatively difficult to distinguish from human written texts compared to other similar models in other benchmarks [^1] https://raid-bench.xyz/ as well as our own datasets. The samples range from 30 to 25K words in length with an average length of around 600 words.
@@ 3.2 Dataset Creation
GPT-4o was used through a Azure OpenAI endpoint [^2] https://azure.microsoft.com/en-us/products/ai-services/openai-service/. command-r-plus and aya-23 were used through cohere’s API platform [^3] https://dashboard.cohere.com/. Rest of the models were used through open router’s [^4] https://openrouter.ai/models API. The Rewritten samples were created by providing the generator LLM with the original text and a random prompt among writing an alternate version, a later update of what happened or a rephrased version of the same text. The samples which returned the exact text or a very similar text were once again regenerated. The partially machine generated texts were created by splitting the text at random locations and the generator was asked to finish the text. The split locations were chosen randomly starting from the 30th word to end of text. This was done to provide the LLM with enough context to better work towards text completion.
@@ 4 Our System
We have experimented with various multilingual transformer models (He et al., 2023), (Conneau, 2019), (Beltagy et al., 2020) with/without additional LSTM (Hochreiter, 1997) or CRF layers (Zheng et al., 2015) through a binary token-classification approach. We found that using additional CRF layer produced better results compared to other setups with the same model. All of the transformer models tested have produced nearly identical results over our test set. However, xlm-longformer gave better results over unseen domains and generators’ texts, and was used in the end given the longer default context length of 16384. The token level predictions by the models were then mapped into word-level predictions. We use the model’s predictions to separate text portions based on perceived authorship.
@@ 5.2 Unseen Domains & Unseen Generators
Table 2: Word level Metrics over Mgtd-bench (.2f)
@@ 5.3 Unseen Domains & Unseen Generators & Non-Native Speakers
Table 3: Overall Metrics over ETS essays (.1f)
@@ 5.4 Unseen Domains & Partially Seen Generators & Adversarial Inputs
We have also tested over raid-bench (Dugan et al., 2025) which consists of texts from 11 generators and 8 domains. among them roughly 10% would be from a seen domain (news articles) while the rest are unseen by our models. The dataset’s texts were also created using various sampling strategies (greedy, random, etc.). The texts were also modified to have adversarial methods including homo-glyphs, mis-spellings, alternative spellings, article deletion etc. Among the 11 generators used, Gpt-4 is one which is similar to the generator whose outputs our model has been trained on (Gpt-4o). However, both of them have different linguistic and stylistic features, similar to how Gpt-4 is different from Gpt-3. We have tested our model’s performance once again upon being trained on our own full training data. Additionally, we have also performed an error analysis to find out what domains, models, attack strategies and decoding strategies effected the model’s performance and to what extent. This can be seen in Figure 2, Figure 3, Figure 4 and Figure 5. The texts were classified as machine generated if at least one third of the tokens within the model’s context length were classified as machine generated. The F1 score obtained with the initial model trained on a single generator was 0.63 and the F1 score grew to 0.79 upon being trained on our full dataset. Evaluation was done directly without performing any preprocessing of the texts and neither were our models trained on texts with any of thse adversarial methods.
Figure 2: F1 scores VS text sampling method used
Figure 3: F1 scores VS the generator’s texts
Figure 4: F1 scores VS each domain’s texts
Figure 5: F1 scores VS adversarial method used in the input texts
@@ 6 Other Observations
Table 4: Median length (words) of original & newly generated parts of the sentences : Language wise
@@ Appendix C Reproducibility
We used multilingual longformer  [^5] https://huggingface.co/hyperonym/xlm-roberta-longformer-base-16384 with an additional CRF layer. The hyper-parameters used for training the models can be seen in Table 9. We built a separate model for each language, the training was done over A100 SXM over 10h each.
@@ Appendix D Other Metrics
The metrics over each type of text for each language and LLM separately can be seen in Table 13, Table 14, Table 15, Table 16, Table 17, Table 18, Table 19, Table 20, Table 21, Table 22, Table 23, Table 24.
@@ Appendix E License
The xlm-longformer model we used was available with an mit license, we are releasing the models and datasets through CC BY-NC 4.0666https://creativecommons.org/licenses/by-nc/4.0/deed.en which permits usage for research purposes.

### v2

@@ 1 Introduction
Source Dataset/Benchmark Samples Languages Generators
(Lee et al., 2022) CoAuthor 1,445 1 1
(Zhang et al., 2024) MixSet 3,600 1 12
(Dugan et al., 2022) RoFT 21,646 1 5
(Macko et al., 2024b) MultiTude 4,070 11 8
(Artemova et al., 2025) Beemo 19,600 1 10
Our Work placeholder 2,447,221 23 12
Table 1: Comparison with other Human-LLM co-authored datasets & benchmarks
@@ 2 Related Works
A major portion of current research in detecting machine-generated content focuses on longer-form writing through binary classification. However, AI-generated misinformation is more likely to cause harm than its use in academia, making the distinction between AI and human-generated texts on social media platforms a critical challenge. Existing methods often struggle with accurately identifying AI-generated content over shorter texts. Moreover, binary classification approaches, which categorize texts as either human or AI-generated (Wang et al., 2024a), (Wang et al., 2024b), (Bhattacharjee et al., 2023), (Zellers et al., 2019b), (Macko et al., 2023), (Ghosal et al., 2023), (Dugan et al., 2024) are less practical in settings where texts could be co-authored by both humans and LLMs. In contrast, binary classification may be more effective for shorter texts commonly found on reviews and social media platforms (Macko et al., 2024a), (Ignat et al., 2024), where content typically consists of one or two sentences. Additionally, some detection works rely on detecting watermarks from AI-generated texts, (Chang et al., 2024), (Dathathri et al., 2024), (Sadasivan et al., 2024), (Zhao et al., 2023) but not all generators utilize watermarking limiting the applicability of such approaches. Few other approaches utilize statistical methods (Mitchell et al., 2023), (Kumarage et al., 2023), (Gehrmann et al., 2019b), (Hans et al., 2024), (Bao et al., 2023), but they can be prone to mis-classification against adversarial methods like rephrasing and humanizing. (Abassy et al., 2024) introduced a 4-way classification as entirely human authored, entirely llm authored, human-edited and llm-authored or llm-edited human-authored. An ideal detection system should be capable of identifying AI-generated content from any generator without depending on watermarking, especially since watermarking techniques may not be effective for shorter texts. Further an ideal detector should be robust against adversarial methods. To properly deal with co-authored text cases, a token classification approach to detect boundaries (Dugan et al., 2022), (Macko et al., 2024b) between machine authored and human authored portions might be more appropriate. Further in cases of AI usage in scenarios like academic cases, users are likely to use a proprietary LLM with better knowledge cutoffs than an open source LLM. Similarly, for AI misuse over social platforms, users are more likely to use a open-sourced model due to better flexibility and privacy. Hence, building models and benchmarks with a appropriate set of LLMs might be more applicable for practical scenarios. Many proprietary systems struggle at the task of fine-grained detection, further a large enough dataset to cover all POS-tag bi-grams of the text boundaries is required for such fine-grained detectors to work well (Kadiyala, 2024). Previous works in the similar direction include (Lee et al., 2022), (Zhang et al., 2024), (Dugan et al., 2023), (Macko et al., 2024b), (Liang et al., 2024) which utilize a dataset of limited size and limited number of generators or those less likely to be used, which might not be enough for a detector to work well on unseen domains and generators’ texts. Further, the task of detection of such human-llm co-authored texts is a harder task compared to binary classification of texts based on authorship (Geng and Trotta, 2025) (Huang et al., 2025).
@@ 3 Dataset
Our dataset consists of around 2.45M samples. We used 12 different LLMs out of which 9 are popular proprietary LLMs : GPT-o1 (OpenAI, 2024), GPT-4o (etal., 2024), Gemini-1.5-Pro (DeepMind, 2024), Gemini-1.5-Flash, Claude-3.5-Sonnet (Anthropic, 2023), Claude-3.5-Haiku, Perplexity-Sonar-Large (Perplexity, 2023), Amazon-Nova-Pro (Intelligence, 2024), Amazon-Nova-Lite. We also included 3 open-source LLMs i.e Aya-23 (Aryabumi et al., 2024), Command-R-Plus (Cohere For AI, 2024), Mistral-large-2411 (Mistral AI, 2024) which produced outputs that are relatively difficult to distinguish from human written texts compared to other similar models in other benchmarks [^1] https://raid-bench.xyz/ as well as our own datasets. The samples range from 30 to 25K words in length with an average length of around 600 words. Table 1 provides an overview of comparison with other datasets and benchmarks for fine-grained detection of human-llm co-authored texts. Our dataset utilizes a better choice of generators which are more likely to be used in practical scenarios where as prior datasets are limited to smaller and limited number of models. Our dataset also compromises of over 100 times more samples, aiding other researchers in the field.
@@ 3.2 Dataset Creation
GPT-4o was used through a Azure OpenAI endpoint [^2] https://azure.microsoft.com/en-us/products/ai-services/openai-service/. Command-r-plus and Aya-23 were used through Cohere’s API platform [^3] https://dashboard.cohere.com/. Rest of the models were used through open router’s [^4] https://openrouter.ai/models API. The Rewritten samples were created by providing the generator LLM with the original text and a random prompt among writing an alternate version, a later update of what happened or a rephrased version of the same text. The samples which returned the exact text or a very similar text were once again regenerated. The partially machine generated texts were created by splitting the text at random locations and the generator was asked to finish the text. The split locations were chosen randomly from between the 30th word up to end of text. This was done to provide the LLM with enough context to better work towards text completion.
@@ 4 Our System
We have experimented with various multilingual transformer models (He et al., 2023), (Conneau, 2019), (Beltagy et al., 2020) with/without additional LSTM (Hochreiter, 1997) or CRF layers (Zheng et al., 2015) through a binary token-classification approach. We found that using additional CRF layer produced better results compared to other setups with the same model. All of the transformer models tested have produced nearly identical results over our test set. However, XLM-longformer gave better results over unseen domains and generators’ texts, and was used in the end given the longer default context length of 16384. The token level predictions by the models were then mapped into word-level predictions. We use the model’s predictions to separate text portions based on perceived authorship. Improving the performance required balancing pre and post boundary POS tag bi-grams to reduce error rates (Kadiyala, 2024).
@@ 5.2 Unseen Domains & Unseen Generators
Table 3: Word level Metrics over Mgtd-bench (.2f) through our models (zero-shot, unseen domains, unseen generators)
@@ 5.3 Unseen Domains & Unseen Generators & Non-Native Speakers
Arabic-Best 96.1 96.1 95.0 95.5
English-Best 99.3 99.0 99.2 99.1
Table 4: Overall Metrics over ETS essays (.1f) through our detectors (zero-shot, unseen generators, unseen domain) VS best submissions (fine-tuned on same generators and domain)
@@ 5.4 Unseen Domains & Partially Seen Generators & Adversarial Inputs
We have also tested over raid-bench (Dugan et al., 2025) which consists of texts from 11 generators and 8 domains. among them roughly 10% would be from a seen domain (news articles) while the rest are unseen by our models. The dataset’s texts were also created using various sampling strategies (greedy, random, etc.). The texts were also modified to have adversarial methods including homo-glyphs, mis-spellings, alternative spellings, article deletion etc. Among the 11 generators used, Gpt-4 is one which is similar to the generator whose outputs our model has been trained on (Gpt-4o). However, both of them have different linguistic and stylistic features, similar to how Gpt-4 is different from Gpt-3. We have tested our model’s performance once again upon being trained on our own full training data. Additionally, we have also performed an error analysis to find out what domains, models, attack strategies and decoding strategies effected the model’s performance and to what extent. This can be seen in Figure 2, Figure 3, Figure 4 and Figure 5. The texts were classified as machine generated if at least one third of the tokens within the model’s context length were classified as machine generated. The F1 score obtained with the initial model trained on a single generator was 0.63 and the F1 score grew to 0.79 upon being trained on our full dataset. Evaluation was done directly without performing any preprocessing of the texts and neither were our models trained on texts with any of the adversarial methods.
Figure 2: F1 scores VS text sampling method used : Sampling strategy did not effect detection capability
Figure 3: F1 scores VS the generator’s texts : instruct/chat variants were harder to detect.
Figure 4: F1 scores VS each domain’s texts : news i.e the only training data domain was easier to detect.
Figure 5: F1 scores VS adversarial method used in the input texts : homo-glyphs are the only real issue, and paraphrases to a small extent, while the rest can be handled through pre-processing.
5.5 Comparisons with proprietary systems
While many proprietary systems claim to have excellent results, they often struggle with unseen domains and generators. Further many systems like ZeroGPT [^5] https://github.com/zerogpt-net/zerogpt-api and GPTZero [^6] https://gptzero.stoplight.io/docs/gptzero-api/ do not provide fine-grained predictions through their API. This would require manually evaluating the results through the UI by counting correct and incorrectly classified word counts. Hence, these systems were not used for comparisons for fine-grained scores. Several users have tried this manfully over a subset (Kadiyala, 2024) only to find a large gap in performance. For binary classifications as well, they have a minimum length threshold to even work, which would become incompatible over a large number of samples of the benchmark.
@@ 6 Other Observations
Table 5: Median length (words) of original & newly generated parts of the sentences - Language wise : Models tend to finish off current sentence after authorship switch quickly before continuing with the rest of the text.
@@ 7.3 Ideal Usage
Acknowledgements
Several stages of this work were supported by grants from Traversaal.ai, Cohere Labs & Microsoft Azure.
@@ Appendix C Reproducibility
We used multilingual longformer  [^7] https://huggingface.co/hyperonym/xlm-roberta-longformer-base-16384 with an additional CRF layer. The hyper-parameters used for training the models can be seen in Table 10. We built a separate model for each language, the training was done over H100 SXM over 10h each.
@@ Appendix D Other Metrics
The metrics over each type of text for each language and LLM separately can be seen in Table 14 to Table 25 respectively..
@@ Appendix E License
The xlm-longformer model we used was available with an mit license, we are releasing the models and datasets through CC BY-NC 4.0888https://creativecommons.org/licenses/by-nc/4.0/deed.en which permits usage for research purposes.